OpenStudy (anonymous):

How do i determine a formula for the polynomial function whose graph has zeros (x+5) ( x +1) (x-4) and a y intercept of -40

OpenStudy (anonymous):

it looks like you have it already. only one step left

OpenStudy (anonymous):

i assume you mean the zeros are -5,-1 and 4

OpenStudy (anonymous):

i tried f(x) = -40 (x+5).... etc yes those are the zeros

OpenStudy (anonymous):

the only thing you don't have is the leading coefficient so write \[p(x)=a(x+5) ( x +1) (x-4) \] then you know \[p(0)=-40\] so write \[a(5\times 1\times -4)=-40\] and solve for \(a\)

OpenStudy (anonymous):

a= 2

OpenStudy (anonymous):

clear or no? y intercept is what you get when \(x=0\) yes, it is 2

OpenStudy (anonymous):

how do i write that out

OpenStudy (anonymous):

2x^3 + 5x^2 + x

OpenStudy (anonymous):

not sure what you mean. you have it

OpenStudy (anonymous):

oh you mean how to you multiply out? you just have to grind it til you find it. do a bunch of multiplications, or else cheat

OpenStudy (anonymous):

ohhh ok .. 2(x+5)(x+1)(x-4) then yeah.. from there i guess i need to do the math.. thats no prob though.. thanks a lot

OpenStudy (anonymous):

yes, as they say "do the math" i like cheating, that way i know it is right

OpenStudy (anonymous):

yes that looks like my graph too.. haha. thank you

OpenStudy (anonymous):

yw